OpenStudy (anonymous):

the equation of the line tangent to y=2secx at ax=pi/4 is ? and i get 2sqr2x-y=squr2/2pi-2sqt2, am i correct??

hartnn (hartnn):

you mean this ? \(2\sqrt 2x -y = (\sqrt 2/2)\pi - 2\sqrt 2\) then yes, its correct :)